At many colleges, the number of full-time-equivalent students, f, is given by
f=n/15
where n is the total number of credits for which students enroll in a given semester. Determine the number of full-time-equivalent students on a campus in which students registered for a total of 25,350 credits.

Answer by user_dude2008(1862)

f = n/15 = 25530/15 = 1702

f = 1702